# ZeroG: Investigating Cross-dataset Zero-shot Transferability in Graphs

### v1

@@ Abstract.
With the development of foundation models such as large language models, zero-shot transfer learning has become increasingly significant. This is highlighted by the generative capabilities of NLP models like GPT-4, and the retrieval-based approaches of CV models like CLIP, both of which effectively bridge the gap between seen and unseen data. In the realm of graph learning, the continuous emergence of new graphs and the challenges of human labeling also amplify the necessity for zero-shot transfer learning, driving the exploration of approaches that can generalize across diverse graph data without necessitating dataset-specific and label-specific fine-tuning. In this study, we extend such paradigms to Zero-shot transferability in Graphs by introducing ZeroG, a new framework tailored to enable cross-dataset generalization. Addressing the inherent challenges such as feature misalignment, mismatched label spaces, and negative transfer, we leverage a language model to encode both node attributes and class semantics, ensuring consistent feature dimensions across datasets. We also propose a prompt-based subgraph sampling module that enriches the semantic information and structure information of extracted subgraphs using prompting nodes and neighborhood aggregation, respectively. We further adopt a lightweight fine-tuning strategy that reduces the risk of overfitting and maintains the zero-shot learning efficacy of the language model. The results underscore the effectiveness of our model in achieving significant cross-dataset zero-shot transferability, opening pathways for the development of graph foundation models. Especially, ZeroG, as a zero-shot method, can even achieve results comparable to those of semi-supervised learning on Pubmed.
@@
ccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
When encountering a completely new problem, humans typically start to compare and connect it with the knowledge that they are familiar with. The same idea is also applied to Machine Learning (ML). ML methods usually focus on classifying instances whose classes have already been seen in training. However, in practice, many applications necessitate classifying instances whose classes have not been seen previously, requiring models to utilize the knowledge gained before to begin reasoning and problem-solving. For example, a model trained on animal images may have never seen a “zebra” during its training. Yet, by understanding the concept of “stripes” and “horse-like” animals from its training, it can successfully identify a zebra when it encounters one for the first time.
@@
Graphs are prevalent across multiple disciplines with diverse applications (Tang et al., 2022, 2023a; Li et al., 2019, 2022). However, we notice that graph learning faces two major challenges: (1) the emergence of new graphs, which makes it impractical to train graph models like Graph Neural Networks (GNNs) (Kipf and Welling, 2016; Veličković et al., 2017) on each individual graph; (2) the difficulty of human labeling due to the complex and diverse nature of graph-structured data. Therefore, zero-shot learning is worth exploring in graph learning, as it allows graph models to generalize and perform reasoning over graphs that have never been seen before. This is also essential for reaching the goal of graph foundation models that can adapt to different data without extra fine-tuning. Due to the lack of detailed discussion and analysis on this task in prior work, our goal is to investigate zero-shot transferability in graphs for conducting cross-dataset transfers, which involves a pre-training phase with multiple datasets that have non-overlapping classes with the test datasets. According to recent surveys (Liu et al., 2023c; Li et al., 2023a), we consider this setting well-suited for the graph foundation model era, emphasizing broad generalization across different data sources. We offer a comprehensive discussion on how our setting differs with other related tasks in Section 2.
Graph transfer learning usually follows a “pre-training and fine-tuning” strategy, which aims to learn some general knowledge for the graph model with easily accessible information to reduce the annotation costs of new graphs. Effective pre-training techniques include node-level comparison (Zhu et al., 2021) or reconstruction (Cheng et al., 2023; Hou et al., 2022), edge-level pretext like edge prediction (Jin et al., 2020), and graph-level contrastive learning such as GraphCL (You et al., 2020). However, to perform zero-shot transfer across datasets, the increased number of pre-training datasets and the absence of fine-tuning labels for the target dataset present unique challenges for these transfer learning methods.
@@ 4.3. Upstream Pre-training
One significant concern of pre-training LMs for source dataset compatibility is the resource-intensive nature of conventional full-parameter pre-training, which also carries a substantial risk of overfitting, especially in zero-shot scenarios (Chen et al., 2023b). To circumvent these limitations, we utilize LoRA (Hu et al., 2021), a parameter-efficient pre-training strategy, which ensures suitability for upstream datasets while minimizing overfitting risks and preserving the LM’s zero-shot learning abilities.
@@
Input: Target dataset $\mathcal{D}$ with $N$ classes, trained language model $LM$, neighborhood aggregation rounds $\lambda$;
Output: Predicted labels for nodes in $\mathcal{D}$
1 Procedure ZeroG($\mathcal{D}$, $LM$, $N$, $\lambda$):
2       Create the prompting node $p$ for target dataset $\mathcal{D}$;
3       Design a dataset description as general semantics for $p$;
4       Generate embeddings $\textbf{h}_{n}$ for all nodes $n\in\mathcal{D}$ using $LM$;
5       Generate embeddings $\textbf{h}_{c_{i}}$ for all classes $c_{i}$, $i\in\{1,\ldots,N\}$ using $LM$;
6       Generate embedding $\textbf{h}_{p}$ for prompting node $p$ using $LM$;
7       Add prompting node $p$ to the graph, connecting to all nodes $n\in$ $\mathcal{D}$;
8       for $j=1$ to $\lambda$  do
9             Perform neighborhood aggregation to update $\textbf{h}_{n}$ for all nodes $n\in$ $\mathcal{D}$;
10
11      foreach  $n$in $\mathcal{D}$ do
12             $y^{\prime}\leftarrow\text{argmax}_{i}(\text{sim}(\textbf{h}_{n},\textbf{h}_{c_{i}})\ |\ i\in\{1,\ldots,N\})$;
13             Assign predicted label $y^{\prime}$ to node $n$;
14
15
Algorithm 1 ZeroG: Downstream Inference
@@ 4.4. Downstream Inference
After pre-training on source datasets, we acquire a pre-trained LM enriched with upstream semantics and structure. This LM is designed for zero-shot transferability, allowing direct application to target datasets without any fine-tuning. We try to unify the pre-training and inference stages, enabling effective knowledge transfer as the datasets in the two phases are made more compatible by following a common template. However, it is still important to distinguish different downstream datasets, in order to capture dataset individuality and achieve dataset-specific optimum. We present the whole downstream inference process in Algorithm 1.
@@ 6.2. LM for Graphs
Recently, language models (LMs) have been increasingly utilized in graph-related tasks. Current methodologies can be classified into three categories, distinguished by the role of LMs in graph-related tasks: as the enhancer, predictor, and alignment component (Li et al., 2023a). Specifically, LM-as-enhancer approaches correspond to enhancing the quality of node embeddings with the help of powerful LMs. For example, TAPE (He et al., 2023) utilizes LLMs to enrich initial node embeddings in GNNs with semantic knowledge relevant to the nodes, notably improving embedding quality. WalkLM (Tan et al., 2023) generates roughly meaningful textual sequences by an automated textualization program and fine-tunes an LM using them. In addition, taking LM as predictors, several works (Wang et al., 2023; Zhao et al., 2023; Ye et al., 2023; Fatemi et al., 2023) utilize LMs to directly make predictions for a wide range of graph tasks, such as classifications and reasonings, within a unified generative paradigm. Lastly, GNN-LM alignment ensures that each encoder’s unique functionalities are preserved while coordinating their embedding spaces at a specific stage. For example, MoMu (Su et al., 2022), MoleculeSTM (Liu et al., 2023b), ConGraT (Brannon et al., 2023), and RLMRec (Wei et al., 2023) merge GNN and LM embeddings, enriching graphs with textual knowledge to boost reasoning. In ZeroG, we follow the approach of using LM as an enhancer, employing the LM for unified graph representation learning to bridge the semantic gap between feature and label spaces across source and target graphs.

### v2

@@ Abstract.
With the development of foundation models such as large language models, zero-shot transfer learning has become increasingly significant. This is highlighted by the generative capabilities of NLP models like GPT-4, and the retrieval-based approaches of CV models like CLIP, both of which effectively bridge the gap between seen and unseen data. In the realm of graph learning, the continuous emergence of new graphs and the challenges of human labeling also amplify the necessity for zero-shot transfer learning, driving the exploration of approaches that can generalize across diverse graph data without necessitating dataset-specific and label-specific fine-tuning. In this study, we extend such paradigms to Zero-shot transferability in Graphs by introducing ZeroG, a new framework tailored to enable cross-dataset generalization. Addressing the inherent challenges such as feature misalignment, mismatched label spaces, and negative transfer, we leverage a language model to encode both node attributes and class semantics, ensuring consistent feature dimensions across datasets. We also propose a prompt-based subgraph sampling module that enriches the semantic information and structure information of extracted subgraphs using prompting nodes and neighborhood aggregation, respectively. We further adopt a lightweight fine-tuning strategy that reduces the risk of overfitting and maintains the zero-shot learning efficacy of the language model. The results underscore the effectiveness of our model in achieving significant cross-dataset zero-shot transferability, opening pathways for the development of graph foundation models [^1] Codes and data are available at https://github.com/NineAbyss/ZeroG.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671982isbn: 979-8-4007-0490-1/24/08copyright: acmcopyrightjournalyear: 2024conference: KDD’24: SIGKDD Conference on Knowledge Discovery and Data Mining; August 25-29, 2024; Barcelona, Spainbooktitle: KDD’24: SIGKDD Conference on Knowledge Discovery and Data Mining, August 25-29, 2024, Barcelona, Spainccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
When encountering a completely new problem, humans typically start to compare and connect it with the knowledge that they are familiar with (Chen et al., 2023c). The same idea is also applied to Machine Learning (ML). ML methods usually focus on classifying instances whose classes have already been seen in training. However, in practice, many applications necessitate classifying instances whose classes have not been seen previously, requiring models to utilize the knowledge gained before to begin reasoning and problem-solving. For example, a model trained on animal images may have never seen a “zebra” during its training. Yet, by understanding the concept of “stripes” and “horse-like” animals from its training, it can successfully identify a zebra when it encounters one for the first time.
@@
Graphs are prevalent across multiple disciplines with diverse applications (Tang et al., 2022, 2023a; Li et al., 2019; Zheng et al., 2021; Li et al., 2022a; Zhang et al., 2022; Zhao et al., 2023a; Yang et al., 2022). However, we notice that graph learning faces two major challenges: (1) the emergence of new graphs, which makes it impractical to train graph models like Graph Neural Networks (GNNs) (Kipf and Welling, 2016; Veličković et al., 2017) on each individual graph; (2) the difficulty of human labeling due to the complex and diverse nature of graph-structured data. Therefore, zero-shot learning is worth exploring in graph learning, as it allows graph models to generalize and perform reasoning over graphs that have never been seen before. This is also essential for reaching the goal of graph foundation models that can adapt to different data without extra fine-tuning. Due to the lack of detailed discussion and analysis on this task in prior work, our goal is to investigate zero-shot transferability in graphs for conducting cross-dataset transfers, which involves a pre-training phase with multiple datasets that have non-overlapping classes with the test datasets. According to recent surveys (Liu et al., 2023c; Li et al., 2023a), we consider this setting well-suited for the graph foundation model era, emphasizing broad generalization across different data sources. We offer a comprehensive discussion on how our setting differs with other related tasks in Section 2.
Graph transfer learning usually follows a “pre-training and fine-tuning” strategy, which aims to learn some general knowledge for the graph model with easily accessible information to reduce the annotation costs of new graphs (Zhao et al., 2024). Effective pre-training techniques include node-level comparison (Zhu et al., 2021) or reconstruction (Cheng et al., 2023; Hou et al., 2022), edge-level pretext like edge prediction (Jin et al., 2020), and graph-level contrastive learning such as GraphCL (You et al., 2020). However, to perform zero-shot transfer across datasets, the increased number of pre-training datasets and the absence of labels for the target dataset present unique challenges for these transfer learning methods.
@@ 4.3. Upstream Pre-training
One significant concern of pre-training LMs for source dataset compatibility is the resource-intensive nature of conventional full-parameter pre-training (Chen et al., 2023d), which also carries a substantial risk of overfitting, especially in zero-shot scenarios (Chen et al., 2023b, 2024b). To circumvent these limitations, we utilize LoRA (Hu et al., 2021), a parameter-efficient pre-training strategy, which ensures suitability for upstream datasets while minimizing overfitting risks and preserving the LM’s zero-shot learning abilities.
@@ 4.4. Downstream Inference
After pre-training on source datasets, we acquire a pre-trained LM enriched with upstream semantics and structure. This LM is designed for zero-shot transferability, allowing direct application to target datasets without any fine-tuning. We try to unify the pre-training and inference stages, enabling effective knowledge transfer as the datasets in the two phases are made more compatible by following a common template. However, it is still important to distinguish different downstream datasets, in order to capture dataset individuality and achieve dataset-specific optimum.
@@ 6.2. LM for Graphs
Recently, language models (LMs) have been increasingly utilized in graph-related tasks (Shen et al., 2021; Shu et al., 2022; Li et al., 2022b). Current methodologies can be classified into three categories, distinguished by the role of LMs in graph-related tasks: as the enhancer, predictor, and alignment component (Li et al., 2023a). Specifically, LM-as-enhancer approaches correspond to enhancing the quality of node embeddings with the help of powerful LMs. For example, TAPE (He et al., 2023) utilizes LLMs to enrich initial node embeddings in GNNs with semantic knowledge relevant to the nodes, notably improving embedding quality. WalkLM (Tan et al., 2023) generates roughly meaningful textual sequences by an automated textualization program and fine-tunes an LM using them. In addition, taking LM as predictors, several works (Wang et al., 2023; Zhao et al., 2023b; Ye et al., 2023; Fatemi et al., 2023; Chen et al., 2024a) utilize LMs to directly make predictions for a wide range of graph tasks, such as classifications and reasonings, within a unified generative paradigm. Lastly, GNN-LM alignment ensures that each encoder’s unique functionalities are preserved while coordinating their embedding spaces at a specific stage. For example, MoMu (Su et al., 2022), MoleculeSTM (Liu et al., 2023b), ConGraT (Brannon et al., 2023), and RLMRec (Wei et al., 2023) merge GNN and LM embeddings, enriching graphs with textual knowledge to boost reasoning. In ZeroG, we follow the approach of using LM as an enhancer, employing the LM for unified graph representation learning to bridge the semantic gap between feature and label spaces across source and target graphs.
@@ 7. Conclusion and Future Work
Acknowledgements.
This work was supported by NSFC Grant No. 62206067, HKUST-HKUST(GZ) 20 for 20 Cross-campus Collaborative Research Scheme C019 and Guangzhou-HKUST(GZ) Joint Funding Scheme 2023A03J0673.